OpenStudy (anonymous):

deshawn and alberto each improved their yards by planting rose bushes and ivy . They bought their supplies from the same store. Deshawn spent $32 dollars on 3 rose bushes and 7 pots of ivy. Alberto spent $30 on 4 rose bushes and 3 pots of ivy. Find the cost of one rose bush and the cost of one pot of ivy.

OpenStudy (anonymous):

32=3r+7i 30=4r+3i

OpenStudy (anonymous):

now set one part of the two equations equal i would recommend multiplying both equations.

OpenStudy (anonymous):

(32=3r+7i)4 (30=4r+3i)3

OpenStudy (anonymous):

(128=12r+28i) (90=12r+9i)

OpenStudy (anonymous):

(128=12r+28i) -(90=12r+9i) 38=19i

OpenStudy (anonymous):

solve for i

OpenStudy (anonymous):

then plug i into one of the original equation and solve for the other variable.

OpenStudy (anonymous):

What about the rose bushes

OpenStudy (anonymous):

i'm not solving the problem for you.

OpenStudy (anonymous):

Well can you at least give me the numbers

OpenStudy (anonymous):

i practically spelled one of the answers for you.

OpenStudy (anonymous):

Well can you help me with the roses than

OpenStudy (anonymous):

i did... then plug i into one of the original equation and solve for the other variable.

OpenStudy (anonymous):

the other variable being r

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

let me know when you get the answer and i'll check it for you.